OpenStudy (anonymous):

Math 30 Trigonometry

OpenStudy (anonymous):

grade twelve math

OpenStudy (anonymous):

but anyways heres the question

OpenStudy (anonymous):

Canadian? :D

OpenStudy (jhannybean):

Did you find a bat signal?

OpenStudy (anonymous):

So it says to determine the general solution for sinx=0 so the solutions are 0 (degrees), 180, and 360 (all in degrees) so I would have three general solution being 360n, 180+360n, and 360+360n Okay so these are correct but they said that it could be combined together to get one solution which would be 180n? howw in the world ?? and yes canadian

OpenStudy (jhannybean):

what is the range for your sine function?

OpenStudy (jhannybean):

I mean interval

OpenStudy (anonymous):

as in period?

OpenStudy (jhannybean):

i.e is it from \((0 \le x \le \dfrac{\pi}{2})\) or is it from \((0 \le x \le 2\pi)\)?

OpenStudy (anonymous):

\[0 \le x \le 360\] assuming

OpenStudy (anonymous):

oh, yeah well it doesnt say in the question

OpenStudy (jhannybean):

\[\sin(x)=0\]\[\sin^{-1}(\sin(x)) = \sin^{-1}(0)\]\[x=\sin^{-1}(0)\]

OpenStudy (jhannybean):

\[\sin(x+2\pi) = 0\]\[\sin^{-1}(\sin(x+2\pi)) = \sin^{-1}(0)\]\[x+2\pi = \sin^{-1}(0)\] I think this would be the proper way of going about solving it.. I'm actually terrible with these. :|

OpenStudy (anonymous):

Mmm I just don't get how the combining of the three works to get 180n...

OpenStudy (kainui):

Here, just make sure you realize what 180n means. It means n is any integer (positive and negative whole numbers both) so try plugging in a few values and we have: \[\Large 180 * 0 =0 \\ \Large 180 * 1 = 180 \\ \Large 180 * 2 = 360 \\ ...\] There are actually infinitely many, and this makes sense, a sine curve continuously wiggles over the x-axis! =D

OpenStudy (anonymous):

 with the graph we can see the general solution is \[xn \pi, n \in Z\] right? So this would give you all your general solutions, and I think your problem maybe asking for specific solutions where then you let n be an integer and plug in numbers.